OpenStudy (anonymous):

what is the equation of the line that passes through the points (2,5) and (1,2)

OpenStudy (anonymous):

First you can calculate out the slope of the line \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] Then using that slope you should be able to determine where the y intercept of that line is. From that information you can put the equation in y=mx+b form.

OpenStudy (anonymous):

what is the answer exactly showing all the work? @whereismymind49

OpenStudy (anonymous):

Well can you first find the slope for me. I'm not going to give you the full answer but I'll help you get to it.

OpenStudy (anonymous):

the full question is What is the equation of the line that passes through the points (2,5) and (1,2)? A. y=3x+5 B. y=3x+2 C. y+3x-5 D. y=3x-1

OpenStudy (anonymous):

@whereismymind49